# Partial Answer of How Transformers Learn Automata

### v1

@@ 2.2 Relation to Attention and Positional Encoding
The attention layer used here acts by taking average (or learned) interactions between the Fourier embeddings, aligning with the fixed or learned frequency interactions. (The frequency stuff is inspired by [2]) Because Fourier transforms are linear, this attention can be realized by fixed-weight projections or by a shallow learned layer.
2.2.1 Proof: Embedding Fourier Composition into Transformer Attention
We aim to realize
$\varphi_{\text{fourier}}(f,g)=F^{-1}(F(f)\cdot F(g))$
using only Transformer components (attention + MLP).
Given token embeddings $f$ and $g$ at each position:
  • Map them into the Fourier basis via fixed sinusoidal projections.
  • For $\mathbb{Z}_{p}$, use $p$-th roots of unity:
Key point: The embedding $\Phi$ is a fixed, non-learned projection.
Within the attention mechanism:
Role Content
Query $Q$ Fourier embedding of $f$: $Q=\Phi(f)$
Key $K$ Fourier embedding of $g$: $K=\Phi(g)$
Value $V$ Fourier embedding of $g$: $V=\Phi(g)$
Standard attention computes:
Instead, we perform:
  • No softmax.
  • No dot product.
  • Elementwise (Hadamard) product in the Fourier domain:
[…]
After Hadamard product:
  • Apply a fixed inverse Fourier transform $F^{-1}$,
  • Recover the composed function in the original domain.
This inverse transform can be implemented as:
  • A fixed linear layer using conjugate Fourier phases, or
  • A small learned complex-valued linear layer.
@@ 3.2 Formal Theorem
The Transformer simulates control ($H$) and memory ($N$) layers separately. Each layer updates in parallel using matrix multiplication (for $H$ action) and addition (for $N$), both of which are associative. Thus, log-depth prefix scan suffices. Attention layers apply the $H$-action to memory embeddings, and MLP layers handle additive accumulation.
@@ 3.4 Answer to the Open Question
Thus, we go beyond Krohn-Rhodes decompositions (which involve only groups) by covering monoid-based structured memory automata.
@@ 3.5 Summary Table
Appendix A Proof of Embedding and attention width
A.1 Embedding Dimension
@@
A.2 Attention Head Width
@@ A.3 Proof of Equality
Appendix B Comparison with the Prime Decomposition and KR Method

### v2

@@ 2.2 Relation to Attention and Positional Encoding
The attention layer used here acts by taking average (or learned) interactions between the Fourier embeddings, aligning with the fixed or learned frequency interactions. (The frequency stuff is inspired by [2]) Because Fourier transforms are linear, this attention can be realized by fixed-weight projections or by a shallow learned layer. The proof of how to embedding is in A.1
@@ 2.4 Proof of Bound of Simulation Automata
3 Generalization to Semigroup and Group Automata
3.1 From Cyclic Groups to Non-Commutative Structures
The key reason the Fourier Transformer simulated modular counters over $\mathbb{Z}_{p}$ was the availability of a Fourier basis: a set of multiplicative characters $\chi_{j}:\mathbb{Z}_{p}\to\mathbb{C}^{\times}$ that formed an orthogonal basis for functions on $\mathbb{Z}_{p}$. Updates like $x\mapsto x+1\bmod p$ became phase rotations in frequency space, implemented as Hadamard multiplications.
To generalize beyond abelian groups, we move to the broader setting of finite semigroups and non-abelian groups—possibly unsolvable or without inverses. Our objective is to simulate automata whose state evolution is governed by algebraic multiplication in such structures, using constant-depth Transformers.
3.2 Fourier Analysis on Semigroups and Groups
Let $S$ be a finite semigroup. The space of functions $f:S\to\mathbb{C}$ admits a representation-theoretic analogue of the Fourier transform.
Let $\mathrm{Rep}(S)=\{\rho^{(1)},\dots,\rho^{(k)}\}$ be a complete set of inequivalent irreducible (complex) matrix representations of $S$, where each
$\rho^{(j)}:S\to\mathbb{C}^{d_{j}\times d_{j}}$
is a homomorphism of semigroups (not necessarily unitary). Then, for any $f:S\to\mathbb{C}$, we define its Fourier transform at $\rho^{(j)}$ as:
$\widehat{f}(\rho^{(j)}):=\sum_{x\in S}f(x)\,\rho^{(j)}(x).$
This maps scalar functions to a tuple of matrices, encoding how $f$ aligns with the algebraic structure of $S$.
If $f,g:S\to\mathbb{C}$, then their semigroup convolution is defined as:
$(f*g)(x):=\sum_{y\in S}f(y)\,g(y^{-1}x),$
where inverses may be generalized using Green’s relations or regular semigroup theory.
When $S$ is a group, this simplifies to the standard group convolution, and Fourier convolution becomes matrix multiplication:
$\widehat{f*g}(\rho^{(j)})=\widehat{f}(\rho^{(j)})\cdot\widehat{g}(\rho^{(j)}).$
3.3 Embedding via Direct Sum of Representations
We embed each element $x\in S$ as:
$\Phi(x):=\bigoplus_{j=1}^{k}\rho^{(j)}(x)\in\mathbb{C}^{D},\quad\text{where }D=\sum_{j}d_{j}^{2}.$
This block-structured vector plays the role of a frequency-domain representation.
Each input token stores this embedding; Transformer layers then simulate updates by applying representation-level matrix multiplication, e.g.:
$\Phi(x\cdot t_{\sigma})=\bigoplus_{j}\rho^{(j)}(x)\,\rho^{(j)}(t_{\sigma}).$
3.4 Definition of Semigroup Automata
Definition 2 (Semigroup Automaton).
Let $S$ be a finite semigroup and $\Sigma$ an input alphabet. A semigroup automaton is a tuple $A=(S,\Sigma,\delta)$ where:
  • The state space is $S$,
  • The transition function is $\delta(s,\sigma):=s\cdot t_{\sigma}$ for each fixed $t_{\sigma}\in S$,
  • Transitions are performed by left-multiplication in $S$.
We now show how such automata can be simulated by shallow Transformers.
3.5 Depth and Width Bounds for Simulation
Theorem 2 (Constant-depth Transformer simulation of semigroup automata).
Let $A=(S,\Sigma,\delta)$ be a semigroup automaton. Let $\mathrm{Rep}(S)=\{\rho^{(1)},\dots,\rho^{(k)}\}$ be a complete set of irreducible complex representations, with total dimension
$D:=\sum_{j=1}^{k}d_{j}^{2}.$
Then there exists a Transformer network with:
  1. 1.
    Depth: $O(1)$, independent of input length $T$,
  2. 2.
    Width: $O(D)$, determined solely by the representation theory of $S$,
which exactly simulates the state evolution of $A$ on any sequence from $\Sigma^{T}$.
Proof.
We construct the Transformer as follows:
  1. 1.
    Embedding. Each state $s\in S$ is encoded as:
    $\Phi(s):=\bigoplus_{j}\rho^{(j)}(s)\in\mathbb{C}^{D}.$
  2. 2.
    Transition. For each input symbol $\sigma$, the fixed right-multiplier $t_{\sigma}\in S$ has representation matrices $\rho^{(j)}(t_{\sigma})$. The transition $s\mapsto s\cdot t_{\sigma}$ becomes:
    $\rho^{(j)}(s\cdot t_{\sigma})=\rho^{(j)}(s)\cdot\rho^{(j)}(t_{\sigma}),$
    which is blockwise matrix multiplication.
  3. 3.
    Locality and Depth. All updates are local (token-wise), and involve only constant-time operations:
    • Each Transformer block performs a fixed blockwise linear transformation,
    • No recurrence or memory across tokens is required,
    • Therefore, total depth is $O(1)$.
  4. 4.
    Width. The embedding dimension is:
    $\text{Width}=D=\sum_{j}d_{j}^{2},$
    where $d_{j}$ is the dimension of $\rho^{(j)}$. This depends only on $S$ and is constant across sequence length.
3.6 Special Case: Group Automata
Definition 3 (Group Automaton).
Let $G$ be a finite group and $\Sigma$ an input alphabet. A group automaton is a semigroup automaton $A=(G,\Sigma,\delta)$ where $G$ is a group and $\delta(g,\sigma)=g\cdot t_{\sigma}$ for $t_{\sigma}\in G$.
Let $\mathrm{Irr}(G)=\{\rho^{(j)}\}$ denote the irreducible unitary representations of $G$. Then:
$\sum_{j}d_{j}^{2}=|G|,$
so the total embedding dimension is $D=|G|$.
Theorem 3 (Transformer simulation of group automata).
Let $A=(G,\Sigma,\delta)$ be a group automaton. Then there exists a Transformer with:
  1. 1.
    Depth: $O(1)$,
  2. 2.
    Width: $O(|G|)$,
which exactly simulates $A$ over any input sequence of length $T$.
3.7 Complexity for Unsolvable Groups
Unsolvable groups (such as $A_{5}$) have high-dimensional irreducible representations and no abelian normal series. For such groups:
  1. 1.
    $\mathrm{Irr}(G)$ remains complete and orthonormal,
  2. 2.
    Representation dimensions $d_{j}$ may be large (e.g., $A_{5}$ has a 5-dimensional irrep),
  3. 3.
    Width $D=\sum_{j}d_{j}^{2}$ grows with group complexity,
  4. 4.
    However, the Transformer simulation remains of constant depth.
Thus, even automata over unsolvable groups are simulable in constant depth — at the cost of increased embedding width and per-token parameter count.
@@ 4.2 Formal Theorem
The Transformer simulates control ($H$) and memory ($N$) layers separately. Each layer updates in parallel using matrix multiplication (for $H$ action) and addition (for $N$), both of which are associative. Thus, log-depth prefix scan suffices. Attention layers apply the $H$-action to memory embeddings, and MLP layers handle additive accumulation. Detail in B.1
@@ 4.4 Answer to the Open Question
Thus, we go beyond Krohn-Rhodes decompositions (which involve only groups) by covering monoid-based structured memory automata.More comparison can be found in C.
@@ 4.5 Summary Table
Appendix A Fourier Module
A.1 Proof: Embedding Fourier Composition into Transformer Attention
We aim to realize
$\varphi_{\text{fourier}}(f,g)=F^{-1}(F(f)\cdot F(g))$
using only Transformer components (attention + MLP).
Given token embeddings $f$ and $g$ at each position:
  • Map them into the Fourier basis via fixed sinusoidal projections.
  • For $\mathbb{Z}_{p}$, use $p$-th roots of unity:
Key point: The embedding $\Phi$ is a fixed, non-learned projection.
Within the attention mechanism:
Role Content
Query $Q$ Fourier embedding of $f$: $Q=\Phi(f)$
Key $K$ Fourier embedding of $g$: $K=\Phi(g)$
Value $V$ Fourier embedding of $g$: $V=\Phi(g)$
Standard attention computes:
Instead, we perform:
  • No softmax.
  • No dot product.
  • Elementwise (Hadamard) product in the Fourier domain:
[…]
After Hadamard product:
  • Apply a fixed inverse Fourier transform $F^{-1}$,
  • Recover the composed function in the original domain.
This inverse transform can be implemented as:
  • A fixed linear layer using conjugate Fourier phases, or
  • A small learned complex-valued linear layer.
Appendix B Representation-theoretic(Semi-direct) Approach
B.1 Proof of Embedding and Attention width
B.1.1 Embedding Dimension
@@
B.1.2 Attention Head Width
@@ B.1.3 Proof of Equality
Appendix C Comparison with the Prime Decomposition and KR Method